OpenStudy (anonymous):

If an object on Planet X weighs 167 N and has a mass of 100 kg, what is the value of gravity, g, for Planet X? A.) 0.167 m/s2 B.) 16.70 m/s2 C.) 1.67 kg D.) 1.67 N*kg E.) 1.67 m/s2

OpenStudy (anonymous):

well you can right off the bat knock off C and D because the unites for acceleration are m/s2. do you know the formula for weight?

OpenStudy (anonymous):

the formula is W=M*G, so if you know W, and you know M, just plug in those values and solve for G, so your final formula will look like: G=W/M.

OpenStudy (anonymous):

the answer will be (E) , because g= 167/ 100

OpenStudy (anonymous):

That is correct, the answer will be 1.67 m/s2